Consider a community consisting of 20 households who depend on a common well for its supply of water. Clean-up and chemical treatment is required to ensure safety of water from the well. The marginal benefits to each household from clean-up and treatment is given by the equation,
MBx=100-4X
If the marginal clean-up and treatment cost is 100 + 20X where X represents the number of clean-ups and chemical treatment per year,
1. Determine the social optimal number of clean-ups and chemical treatment per year.
2. Compute the net social welfare benefit to each household if the households share the cost of clean-ups and treatment
3. Compute the total social welfare benefit to each household if half of the households pay half the cost of clean-ups and treatment
4. Compute the amount of free riding from the two groups in 3 above
1
Expert's answer
2020-04-20T10:29:41-0400

1."20\times(100-4x)=100+20x"

2000-80x=100+20x

1900=100x

x=19

2."20\times(100-4\times19)-(100+20\times19)=2000-76-100-380=1444"

"\frac{1444}{20}=72.2"

3."100+20\times19=480"

"\frac{480}{2}=240"

4."6\times(100-4x)=0"

600-24x=0

x=25